Grade 12Physical Chemistry

The pKa of a weak acid (HA) is 4.5. The pOH of an aqueous buffer solution of HA in which 50% of the acid is ionized is ?

sudhanshu
12 Years agoGrade 12

2 Answers

ajju
9 Years ago

ph =pka=4.5          ; this is because the ratio log [A^-]/[HA]=log o.5/0.5=0

                                       hence, ph=pka in the formulla ph=pka+log [A^-]/[HA]

hence, poh= 14-ph=14-4.5=9.5    

Rishi Sharma
6 Years ago
Dear Student,
Please find below the solution to your problem.

ph =pka=4.5
this is because the ratio log [A^-]/[HA]=log o.5/0.5=0
hence, ph=pka
in the formulla ph=pka+log [A^-]/[HA]
hence, poh= 14-ph=14-4.5=9.5

Thanks and Regards